OpenStudy (xbeenchantedx):

Solve the equation below. Check for extraneous solutions. cubed root (4x-2) = 2 cubed root (x+6)

OpenStudy (xbeenchantedx):

\[\sqrt[3]{4x-2} = 2\sqrt[3]{x+6}\]

OpenStudy (tkhunny):

Have you considered cubing both sides?

OpenStudy (xbeenchantedx):

@tkhunny I have and I ended up with x=7, but when I plugged it back in as x in the original equation to check for extraneous solutions, I end up with a confusing result.

OpenStudy (xbeenchantedx):

\[\sqrt[3]{4(7)-2} = 2\sqrt[3]{(7)+6} -> \sqrt[3]{26} = 2\sqrt[3]{13}\]

OpenStudy (xbeenchantedx):

I'm not sure if the x I got was right.

OpenStudy (tkhunny):

Show me what you got after cubing both sides. Did you get 2 or 8 on the right-hand side?

OpenStudy (xbeenchantedx):

On my first attempt I got 2. I think I forgot to also cube the 2. On my second attempt I got 8, but my answer ended up as x=22 and when I checked my answer, I ended up with weird results.

OpenStudy (tkhunny):

4x-2 = 8(x+6) 4x-2 = 8x+48 4x-50 = 8x -50 = 4x x = -25/2 -- Try that. Be more careful.

OpenStudy (xbeenchantedx):

I see what I got wrong (I accidentally wrong 42 instead of 48). Can you show me how to check my answer using that answer? I tried and I'm either making mistakes or I don't know how to simplify radicals.

OpenStudy (mathstudent55):

\(\large \sqrt[3]{4x-2} = 2\sqrt[3]{x+6}\) \(\large \sqrt[3]{ 4 \times \left(- \dfrac{25}{2} \right) -2} = 2\sqrt[3]{- \dfrac{25}{2} +6}\)

OpenStudy (xbeenchantedx):

@mathstudent55 When I reduce that, I get \[\sqrt[3]{-52}=2\sqrt[3]{-13/2}\] From there, I am stuck. Is that all I'm supposed to get? Or can I simplify it even more?

OpenStudy (mathstudent55):

\(\large \sqrt[3]{ 4 \times \left(- \dfrac{25}{2} \right) -2} = 2\sqrt[3]{- \dfrac{25}{2} +6}\) \(\large \sqrt[3]{ -50 -2} = 2\sqrt[3]{- \dfrac{25}{2} +\dfrac{12}{2}}\) \(\large \sqrt[3]{ -52} = 2\sqrt[3]{- \dfrac{13}{2} }\) \(\large \sqrt[3]{ -52} = \sqrt[3]{8}\sqrt[3]{- \dfrac{13}{2} }\) \(\large \sqrt[3]{ -52} = \sqrt[3]{8 \times \left(- \dfrac{13}{2} \right)}\) \(\large \sqrt[3]{ -52} = \sqrt[3]{- 52}\)

OpenStudy (mathstudent55):

The solution works, so it's not extraneous.

OpenStudy (xbeenchantedx):

Ohhh now I get it. Thank you so much.

OpenStudy (tkhunny):

It's also in the Domain, so it's not extraneous.